# ShadowNav: Autonomous Global Localization for Lunar Navigation in Darkness

### v2

@@ Abstract
The ability to determine the pose of a rover in an inertial frame autonomously is a crucial capability necessary for the next generation of surface rover missions on other planetary bodies. Currently, most on-going rover missions utilize ground-in-the-loop interventions to manually correct for drift in the pose estimate and this human supervision bottlenecks the distance over which rovers can operate autonomously and carry out scientific measurements. In this paper, we present ShadowNav, an autonomous approach for global localization on the Moon with an emphasis on driving in darkness and at nighttime. Our approach uses the leading edge of Lunar craters as landmarks and a particle filtering approach is used to associate detected craters with known ones on an offboard map. We discuss the key design decisions in developing the ShadowNav framework for use with a Lunar rover concept equipped with a stereo camera and an external illumination source. Finally, we demonstrate the efficacy of our proposed approach in both a Lunar simulation environment and on data collected during a field test at Cinder Lakes, Arizona.
@@ 1 Introduction
Space missions that require long-range autonomous driving on the surface of extra planetary bodies have gained significant interest lately especially for the Lunar surface. For example, in the latest Decadal Survey [National Academies of Sciences, Engineering, and Medicine, 2022b], the Endurance-A Lunar rover was recommended to be implemented as a strategic class mission as the highest priority of the Lunar Discovery and Exploration Program. This mission proposal calls for a $2000\text{\,}\mathrm{km}$ traverse in the South Pole-Aitken (SPA) Basin to collect $100\text{\,}\mathrm{kg}$ of samples for delivery to Artemis astronauts. This mission concept study [Keane et al., 2022] identifies several key capabilities required to complete this mission:
@@
Figure 1: The ShadowNav concept relies on using crater rims as landmarks. As the rover begins driving (a), its uncertainty in the global frame increases and must be corrected for. By equipping the Lunar rover with a stereo camera and an external illumination source, the leading edges of detected craters (shown in green) are used to associate the crater with known ones from an offline map and decrease the pose uncertainty. As the rover continues to drive, (b) this process repeats intermittently to be able to continue mission operations.
@@
This level of autonomy and drive distance would be an order of magnitude larger than any previous surface mission on an extra planetary body. For example, as the current state-of-the-art in beyond Earth surface autonomy, Mars 2020 has driven $11.6\text{\,}\mathrm{km}$ autonomously over the course of a year and a half and the longest individual autonomous drive was approximately $700\text{\,}\mathrm{m}$ [Verma et al., 2023a]. A key bottleneck that limits longer autonomous drives is the problem of absolute localization, which is the problem of determining the position and orientation of the vehicle in an inertial frame. Without a reliable means of accomplishing autonomous localization autonomously, the vehicle must stop and wait for ground-in-the-loop operators to manually perform absolute localization. In the case of surface rovers used on other planetary bodies, existing relatively localization methods experience approximately 2% drift and this limits autonomous driving to approximately $500\text{\,}\mathrm{m}$ before a manual absolute localization is triggered once the error exceeds $10\text{\,}\mathrm{m}$  [Maimone et al., 2007, Maimone et al., 2022].
@@
Currently however, the entirety of the Lunar surface does not have continuous communication with Earth and having the vehicle wait until communications with the ground station is possible will severely slow down the traverse rates. The lack of frequent absolute localization for the rover would lead to errors greater than the maximum $10\text{\,}\mathrm{m}$ localization error which would present significant risks to the mission through deviations from the desired trajectory. For proposed missions such as Endurance-A, the requirement to drive longer distances on the order of several kilometers a day necessitates the development of an autonomous and performant technique to perform absolute localization onboard the vehicle and without ground-in-the-loop intervention.
@@
In this work, we draw inspiration from recent techniques that have made use of craters as landmarks for performing absolute localization on the Moon for daytime driving [Matthies et al., 2022, Daftry et al., 2023]. For Lunar rover missions, craters have the potential to serve as landmarks as the average distance between craters of $\geq$$10\text{\,}\mathrm{m}$ in diameter is $100\text{\,}\mathrm{m}$ on terrain with relatively fresh craters and $10\text{\,}\mathrm{m}$ on terrain with old craters [Hiesinger et al., 2012]. Further, the Lunar Reconnaissance Orbiter Camera (LROC) provides digital elevation models (DEMs) with a resolution between $0.5\text{\,}\mathrm{m}$-$5\text{\,}\mathrm{m}$ per pixel [Robinson et al., 2010] and there are some DEMs within Permanently Shadowed Regions (PSRs) [Cisneros et al., 2017]. However, unlike the daytime driving case, the lack of natural light available when driving within a PSR or during the Lunar night limits what can be used as a landmark and the range at which the landmarks can be observed.
@@
To address this challenge of degraded perception conditions, recent mission concepts have proposed equipping rover systems with a stereo camera and an illumination source for driving at night and in the dark [Keane et al., 2022, Robinson and Elliott, 2022, National Academies of Sciences, Engineering, and Medicine, 2022a]. In this work, we propose an autonomous global localization technique for use in such mission architectures that use a stereo camera with an external illumination source in order to detect crater rims in darkness. Global localization is accomplished by matching the detected crater rims against known craters from an orbital image which is done using our novel Q-Score metric for ranking potential crater matches. A particle filter with the the Q-Score metric is utilized to estimate the absolute position amongst the uncertainty and nonlinearity of crater rim matches.
@@
Summary of Contributions: A preliminary version of our proposed approach was presented in [Cauligi et al., 2023]. In this revised and extended version, we provide the following additional contributions:
@@
    Evaluation of optimal trajectories to view a landmark for absolution localization within simulated Lunar environment.
@@
Organization: We begin in Section 2 by reviewing prior work on global localization for surface rover missions and the underlying techniques necessary for detecting craters as landmarks. In Section 3, we present our proposed architecture on using stereo cameras to detect the leading edges of craters and the particle filter in which these craters as used as landmarks for global localization. Section 4 presents the datasets used in our work, including (1) a photorealistic simulation environment used to generate synthetic stereo camera data and (2) nighttime datasets collected during field tests conducted in Cinder Lakes, Arizona. Next, Section 5 provides numerical results using these aforementioned datasets to validate the efficacy of our proposed approach. Finally, we conclude in Section 6 and present directions of future work.
@@ 2 Related Work
Existing Mars rover missions provide the benchmark for state-of-the-art in autonomous surface rover capabilities. While autonomy continues to play an increased role in Mars rover operations [Verma et al., 2023b], global localization is still performed using manual ground-in-the-loop interventions. As the visual odometry used on-board for relative localization accures approximately 2-3% error [Johnson et al., 2008], these interventions are triggered when the error in absolute localization increases beyond an acceptable threshold. For the proposed Endurance-A mission [Keane et al., 2022], the mission concept-of-operations (ConOps) requires that the rover maintain position knowledge in the inertial frame with less than $10\text{\,}\mathrm{m}$ error. Under the assumption of 2% relative localization error for relative localization, the Endurance-A mission then expects to perform an absolute localization update approximately every $300\text{\,}\mathrm{m}$.
@@
In this work, we develop a framework that relies on craters as landmarks for use in absolute localization. Prior work by [Hiesinger et al., 2012] has estimated the frequency of craters of size $10\text{\,}\mathrm{m}$ to be every $100\text{\,}\mathrm{m}$, which allows for craters to be encountered frequently enough for an global localization update to be performed at the required intervals of $300\text{\,}\mathrm{m}$ for Endurance-A. Moreover, we build upon prior work that has demonstrated the efficacy of using craters as landmarks for use in daytime Lunar driving [Matthies et al., 2022].
@@
In order to detect these craters as landmarks in this work, we utilize stereo cameras as the primary sensor. Stereo cameras have been one of the primary perception technologies for rover navigation on the extraplanetary bodies, most notably the Mars rover missions and is included as one of the primary perception sensors in proposed Lunar missions [Keane et al., 2022, Robinson and Elliott, 2022]. In addition to classical methods for crater detection, recent approaches such as Semi-Global Block Matching (SGBM) [Hirschmuller, 2007] and  Domain-Invariant Stereo Matching Networks (DSMNets) [Zhang et al., 2020] have demonstrated promise in using deep learning for crater detection using stereo.
@@
Crater detection is a crucial aspect of Lunar navigation and for utilizing craters as landmarks for accurate absolute localization. Various methodologies have been explored to achieve accurate crater detection in challenging lunar environments. Studies such as Lunarnet [Liounis et al., 2019] have utilized convolutional neural networks (CNNs) to detect craters from orbital images captured by instruments on-board, e.g., the LROC [Robinson et al., 2010] spacecraft. Lunarnet aimed to determine spacecraft pose by analyzing crater patterns, although challenges such as false positives and sensitivity to sunlight were encountered. Algorithmic approaches have also been investigated. In a comparison of crater-detection algorithms, [Woicke et al., 2018] evaluated six different methods based on criteria such as detection rates, false detection, accuracy, robustness and run-time. These methods include edge detection, identifying bright or shaded portions of a crater, template matching, and supervised learning-based techniques. Despite their merits, these algorithms often struggle with complex scenes, overlapping craters, and lighting conditions. Additionally, works such as DeepMoon [Silburt et al., 2019] employ elevation maps generated from radar or lidar data for crater detection. Template matching, as seen in PyCDA [Klear, 2018], offers an open-source solution for automated crater detection using CNNs and per-pixel likelihood estimation. While these methods have contributed to advancements in crater detection, challenges remain in extending capabilities to nighttime or shadowed operations and improving accuracy in diverse lunar terrains.
@@
In order to perform an absolute localization update there have been several proposed methods. In [Hwangbo et al., 2009], the authors propose a localization procedure that matches an observed rover image with an orbital map. However, this results in a deterministic estimate of the robot belief since it neglects the rover motion model. Prior works have also proposed the use of Lunar satellites to provide inertial position updates for Lunar rovers [Bhamidipati et al., 2023, Cortinovis et al., 2024], but current orbiters do not provide this capability and a crater-based localization approach would still fill in temporal gaps between such orbiter updates. [Wu et al., 2019] presents a purely data driven model using a convolutional neural network trained on synthetic data to match the rover observations with orbital imagery. Similar to our approach, [Franchi and Ntagiou, 2022] presents a particle filtering technique which uses a Siamese neural network to compare rover monocular camera imagery with orbital imagery to assign each particle a likelihood weight. The authors in [Daftry et al., 2023] propose a similar approach for Lunar absolute localization known as LunarNav. However, LunarNav focuses on the daytime localization problem and therefore considers different methods of crater matching that rely on greater knowledge of the surface geometry than available in the nighttime case.
@@
A crucial component of localization in a Lunar environment is the problem of crater matching. For instance, LunarNav [Daftry et al., 2023] introduces a novel approach that leverages crater matching for absolute localization for daytime Lunar driving using particle filtering and parametric matchins using an on-board Lidar. Notably, the authors in [Silburt et al., 2019] and [Silvestrini et al., 2022] use neural networks to match craters, however these require a well labelled library to be known in advance for template matching.
@@ 3.1 Image Enhancement
The first step in our system is to utilize Contrast Limited Adaptive Histogram Equalization (CLAHE) [Yadav et al., 2014] to enhance the raw images. In this work, we consider a single light source with consistent illumination provided across the field of light. Therefore, there is a significant decrease in pixel intensity values captured by a camera as the range increases of the physical region being imaged. Since CLAHE is adaptive, it equalizes the image within sub-regions, producing a more consistent image brightness to use as input to stereo algorithms. A potential downside of CLAHEs is that it can magnify noise within the image which is explored within this work.
@@ 3.2 Stereo
The second step within our system is to compute stereo. This is accomplished utilizing [Kogan,] and SGBMs [Hirschmuller, 2007]. The disparity and range maps produced provide the input for crater detection algorithms. In order to handle noisy disparity measurements in the far range, an additional cleanup step is used after stereo is computed.
Stereo Refinement: In order to perform stereo refinement, it is assumed that the terrain will be mostly planar with ranges increasing as the pixel indices get closer to the top row of the image. Therefore, to filter the far range noisy stereo measurements, the average range for all image rows are first computed. Next, these average row values are searched starting at ranges beyond a parameter, $20\text{\,}\mathrm{m}$, in this work. During this search, if the range decreases for a parameterized number of rows consistently, then all computed stereo values after this peak are removed.
@@ 3.3 Crater Leading Edge Detection
After stereo is computed, this output is utilized to perform crater leading edge detection. This is done with two components. The first is to detect discontinuities within the disparity and range images and the second is to connect these discontinuity locations into a single rim via a contour detection.
Discontinuity Detection: To perform the discontinuity detection, vertical columns within the disparity image are searched from bottom to top with holes removed from the search. Along the search if both a disparity discontinuity and range discontinuity are larger than set thresholds then the pixel at the start of the discontinuity is marked within a mask.
@@ 3.4 Absolute Localization via Particle Filter
Here, we provide an overview of the proposed ShadowNav particle filtering approach. First, we provide further details on the Q-Score metric that is used in the belief update step.
3.5 Orbital Crater Ground Truth
In this work, we consider only specific ground truth craters which require a planned trajectory to drive towards. To this end, only the specific landmark craters to be used for global localization are labeled. When the estimated particle filter position was not within $20\text{\,}\mathrm{m}$ of the ground truth rim, the particle filter was not run and the position was propagated by a simulated noisy relative localization estimate update specified in Section 3.6. Secondly, only the front half of a crater is used as ground truth. However, since the landmark crater can be approached from any angle, the entire crater rim is labeled and the front half of the rim is determined dynamically based on the heading of camera at a specific time. Only the front edge of the crater is used as ground truth as our detection algorithm is only attempted to detect the front half and using the entire crater rim for matching could lead to false positives.

### v3

@@ Abstract
The ability to determine the of a rover in an inertial frame autonomously is a crucial capability necessary for the next generation of surface rover missions on other planetary bodies. Currently, most on-going rover missions utilize ground-in-the-loop interventions to manually correct for drift in the estimate and this human supervision bottlenecks the distance over which rovers can operate autonomously and carry out scientific measurements. In this paper, we present ShadowNav, an autonomous approach for global localization on the Moon with an emphasis on driving in darkness and at nighttime. Our approach uses the leading edge of Lunar craters as landmarks and a particle filtering approach to associate detected craters with known ones on an offboard map. We discuss the key design decisions in developing the ShadowNav framework for use with a Lunar rover concept equipped with a stereo camera and an external illumination source. Finally, we demonstrate the efficacy of our proposed approach in both a Lunar simulation environment and on data collected during a field test at Cinder Lakes, Arizona.
{IEEEkeywords}
computer vision, extreme environments, GPS-denied operation, localization, perception
@@ 1 INTRODUCTION
[…]
Space missions that require long-range autonomous driving on the surface of planetary bodies have gained significant interest lately especially for the Lunar surface. For example, in the latest Decadal Survey [1], the Endurance-A Lunar rover was recommended to be implemented as a strategic class mission as the highest priority of the Lunar Discovery and Exploration Program. This mission proposal calls for a $2000\text{\,}\mathrm{km}$ traverse in the South Pole-Aitken (SPA) Basin to collect $100\text{\,}\mathrm{kg}$ of samples for delivery to Artemis astronauts. This mission concept study [2] identifies several key capabilities required to complete this mission:
@@
Figure 1: The ShadowNav concept relies on using crater rims as landmarks. As the rover begins driving (a), its uncertainty in the global frame increases and must be corrected. By equipping the Lunar rover with a stereo camera and an external illumination source, the leading edges of detected craters (shown in green) are used to associate the crater with known ones from an offline map and decrease the pose uncertainty. As the rover continues to drive, (b) this process repeats intermittently to be able to continue mission operations.
@@
This level of autonomy and drive distance would be an order of magnitude larger than any previous surface mission on a planetary body. For example, as the current state-of-the-art in beyond Earth surface autonomy, Mars 2020 has driven $11.6\text{\,}\mathrm{km}$ autonomously over the course of a year and a half and the longest individual autonomous drive was approximately $700\text{\,}\mathrm{m}$ [3]. A key bottleneck that limits longer autonomous drives is the problem of absolute localization which is determining the position and orientation of the vehicle in an inertial frame. Without a reliable means of accomplishing localization autonomously, the vehicle must stop and wait for ground-in-the-loop operators to manually perform absolute localization. In the case of surface rovers used on planetary bodies, existing relatively localization methods experience approximately 2% drift and this limits autonomous driving to approximately $500\text{\,}\mathrm{m}$ before a manual absolute localization is triggered once the error exceeds $10\text{\,}\mathrm{m}$  [4, 5].
@@
Currently however, the entirety of the Lunar surface does not have continuous communication with Earth. Planning for a rover to wait until communications with the ground station is possible will severely slow down the traverse rate. The lack of frequent absolute localization for the rover would lead to errors greater than the maximum $10\text{\,}\mathrm{m}$ localization error of Endurance-A, which could present significant risks to the mission. For proposed missions such as Endurance-A, the requirement to drive longer distances on the order of several kilometers a day necessitates the development of an autonomous and performant technique for absolute localization onboard the vehicle without ground-in-the-loop intervention.
@@
In this work, we draw inspiration from recent techniques that have made use of craters as landmarks for performing absolute localization on the Moon during daytime driving [6, 7]. For Lunar rover missions, craters have the potential to serve as landmarks as the average distance between craters of $\geq$$10\text{\,}\mathrm{m}$ in diameter is $100\text{\,}\mathrm{m}$ on terrain with relatively fresh craters and $10\text{\,}\mathrm{m}$ on terrain with old craters [8]. Further, the Lunar Reconnaissance Orbiter Camera (LROC) provides digital elevation models (DEMs) with a resolution between $0.5\text{\,}\mathrm{m}$-$5\text{\,}\mathrm{m}$ per pixel [9] and there are some DEMs within Permanently Shadowed Regions (PSRs) [10]. However, unlike the daytime driving case, the lack of natural light available when driving within a PSR or during the Lunar night limits what can be used as a landmark and the range at which the landmarks can be observed.
@@
To address this challenge of degraded perception conditions, recent mission concepts have proposed equipping rover systems with a stereo camera and an illumination source for driving at night and PSRs [2, 11, 12]. In this work we propose an autonomous global localization technique that uses a stereo camera with an external illumination source to detect crater rims in darkness for use in such mission architectures. Global localization is accomplished by matching and scoring the detected crater rims against known craters from an orbital image using our novel Q-Score metric. A particle filter is used to estimate the absolute position amongst the uncertainty and nonlinearity of crater rim matches and corresponding Q-Scores.
@@
Summary of Contributions: A preliminary version of our proposed approach was presented in [13]. In this revised and extended version, we provide the following additional contributions:
@@
    Evaluation of optimal trajectories to view a landmark for absolution localization within a simulated Lunar environment.
@@
Organization: We begin in Section 2 by reviewing prior work on global localization for surface rover missions and the underlying techniques necessary for detecting craters as landmarks. In Section 3, we present our proposed architecture on using stereo cameras to detect the leading edges of craters and the particle filter in which these craters are used as landmarks for global localization. Section 4 presents the datasets used in our work, including (1) a photorealistic simulation environment used to generate synthetic stereo camera data and (2) nighttime datasets collected during field tests conducted in Cinder Lakes, Arizona. Next, Section 5 provides numerical results using these aforementioned datasets to validate the efficacy of our proposed approach. Finally, we conclude in Section 6 and present directions of future work.
@@ 2 RELATED WORK
Existing Mars rover missions provide the benchmark for state-of-the-art in autonomous surface rover capabilities. Visual odometry used on-board these Mars rovers for relative localization accrues approximately 2-3% error [14]. For the proposed Endurance-A mission [2], the mission concept-of-operations (ConOps) requires that the rover maintain position knowledge in the inertial frame with less than $10\text{\,}\mathrm{m}$ error Under the assumption of 2% relative localization error, the Endurance-A mission then expects to perform an absolute localization update approximately every $300\text{\,}\mathrm{m}$.
@@
In this work, we develop a framework that relies on craters as landmarks to determine Prior work by [8] has estimated the frequency of craters of size $10\text{\,}\mathrm{m}$ to be every $100\text{\,}\mathrm{m}$, which allows for craters to be encountered frequently enough for a global localization update to be performed at the required intervals of $300\text{\,}\mathrm{m}$ for Endurance-A. Moreover, we build upon prior work that has demonstrated the efficacy of using craters as landmarks for use in daytime Lunar driving [6].
@@
To detect craters as landmarks in this work, we utilize stereo cameras as the primary sensor. Stereo cameras have been one of the primary perception technologies for rover navigation on planetary bodies, most notably the Mars rover missions, and are included as one of the primary perception sensors in proposed Lunar missions [2, 11]. In addition to classical methods for crater detection, recent approaches such as Semi-Global Block Matching (SGBM) [24] and  Domain-Invariant Stereo Matching Networks (DSMNets) [25] have demonstrated promise in using deep learning for crater detection using stereo.
@@
Crater detection is a crucial aspect of Lunar navigation and for using craters as landmarks for accurate absolute localization. Various methodologies have been explored to achieve accurate crater detection in challenging Lunar environments. Studies such as Lunarnet [26] have used convolutional neural networks (CNNs) to detect craters from orbital images captured by instruments on-board, e.g., the LROC [9] spacecraft. Lunarnet aimed to determine spacecraft pose by analyzing crater patterns, although challenges such as false positives and sensitivity to sunlight were encountered. Other algorithmic approaches have also been investigated. In a comparison of crater-detection algorithms, [27] evaluated six different methods based on criteria such as detection rates, false detection, accuracy, robustness, and run-time. These methods include edge detection, identifying bright or shaded portions of a crater, template matching, and supervised learning-based techniques. Despite their merits, these algorithms often struggle with complex scenes, overlapping craters, and lighting conditions. Additionally, works such as DeepMoon [28] employ elevation maps generated from radar or lidar data for crater detection. Template matching, as seen in PyCDA [29], offers an open-source solution for automated crater detection using CNNs and per-pixel likelihood estimation. While these methods have contributed to advancements in crater detection, challenges remain in extending capabilities to nighttime or shadowed operations and improving accuracy in diverse Lunar terrains.
@@
There have been several proposed methods to perform an absolute localization update. In [30], the authors propose a localization procedure that matches an observed rover image with an orbital map. However, this results in a deterministic estimate of the robot belief since it neglects the rover motion model. Prior works have also proposed the use of Lunar satellites to provide inertial position updates for Lunar rovers [31, 32, 33], but current orbiters do not provide this capability and a crater-based localization approach would still fill in temporal gaps between such orbiter updates. [34] presents a purely data driven model using a CNN trained on synthetic data to match the rover observations with orbital imagery. Similar to our approach, [35] presents a particle filtering technique which uses a Siamese neural network to compare rover monocular camera imagery with orbital imagery to assign each particle a likelihood weight. The authors in [7] propose a similar approach for Lunar absolute localization known as LunarNav. However, LunarNav focuses on the daytime localization problem and therefore considers different methods of crater matching that rely on greater knowledge of the surface geometry than available in the nighttime case.
@@
A crucial component of localization in a Lunar environment is the problem of crater matching. For instance, LunarNav [7] introduces a novel approach that leverages crater matching for absolute localization for daytime Lunar driving using particle filtering and parametric matching using an on-board lidar. Notably, the authors in [28] and [36] use neural networks to match craters, however these require a well labelled library to be known in advance for template matching.
@@ 3.2 Image Enhancement
The first step in our system is to utilize Contrast Limited Adaptive Histogram Equalization (CLAHE) [40] to enhance the raw images. In this work, we consider a single light source with consistent illumination provided across the field of light. Therefore, there is a significant decrease in pixel intensity values captured by a camera as the range of the physical region being imaged increases. Since CLAHE is adaptive, it equalizes the image within sub-regions, producing a more consistent image brightness to use as input to stereo algorithms. A potential downside of CLAHE is that it can magnify noise within the image which is explored within this work.
@@ 3.3 Stereo
The second step within our system is to compute stereo. This is accomplished using mrcal [41] and SGBM [24]. The disparity and range maps produced provide the input for crater detection algorithms. An additional cleanup step is used after stereo is computed to handle noisy disparity measurements in the far range.
Stereo Refinement: It is assumed that the as the range of terrain increases, the corresponding pixels will be closer to the top of an image. Therefore, to filter the far range noisy stereo measurements, the average range for all image rows are first computed. Next, these average row values are searched starting at ranges beyond a parameter ($20\text{\,}\mathrm{m}$ in this work). During this search, if the range decreases for a parameterized number of rows consistently, then all computed stereo values after this peak are removed.
@@ 3.4 Crater Leading Edge Detection
After stereo is computed, this output is used to perform crater leading edge detection in two steps. The first is to detect discontinuities within the disparity and range images and the second is to connect these discontinuity locations into a single rim via a contour detection.
Discontinuity Detection: To perform the discontinuity detection, vertical columns within the disparity image are searched from bottom to top with holes removed from the search. Along the search if both a disparity discontinuity and range discontinuity are larger than set thresholds, then the pixel at the start of the discontinuity is marked within a mask.
@@ 3.5 Absolute Localization via Particle Filter
Here, we provide an overview of the proposed ShadowNav particle filtering approach used to handle measurement uncertainties stemming from stereo and disparity computation errors. First, we provide details on Next, we describe our Q-Score metric used in the belief update step. Finally, we present our approach for a filtering system using a particle filter.
3.5.1 Orbital Crater Ground Truth
In this work, we consider only specific ground truth craters which require a planned trajectory to drive towards. To this end, only the specific landmark craters to be used for global localization are labeled. When the estimated particle filter position was not within $20\text{\,}\mathrm{m}$ of the ground truth rim, the particle filter was not run and the position was propagated by a simulated noisy relative localization estimate update specified in Section LABEL:label:rel-loc. Further, only the front half of a crater is used as ground truth. However, since the landmark crater can be approached from any angle, the entire crater rim is labeled and the front half of the rim is determined dynamically based on the heading of camera at a specific time. Because using the entire crater rim for matching could lead to false positives, only the front edge of the crater is used as ground truth as our detection algorithm only attempts to detect the front half.